OpenStudy (anonymous):

How many pounds of walnuts that cost $7.60 per pound must be mixed with 6 pounds of cashews that costs $12 per pound to make a mixture of nuts that costs $9.25 per pound? Select one: a. 8 pounds b. 10.5 pounds c. 10 pounds d. 12 pounds

OpenStudy (aum):

Cost of "x" ponds of walnuts = $7.60x Cost of 6 pounds of cashews = $12 * 6 = $72 Total weight = (x+6) pounds Total cost = $(7.60x + 72) Cost per pound of the mixture = $(7.60x + 72) / (x + 6) = $9.25 Solve for x.

OpenStudy (anonymous):

but how do you find x thats what i dont get

OpenStudy (aum):

(7.60x + 72) / (x + 6) = 9.25 multiply both sides by (x+6) 7.6x + 72 = 9.25(x+6) 7.6x + 72 = 9.25x + 55.5 isolate x. First subtract 7.6x 72 = 1.65x + 55.5 subtract 55.5 16.5 = 1.65x divide by 1.65 10 = x or x = 10 pounds.

OpenStudy (anonymous):

do you divid by 1.65 both sides nd thx

OpenStudy (aum):

Yes. All operations should be done on both sides of the equation to keep the equation still valid.

OpenStudy (anonymous):

oh ok thnxs

OpenStudy (aum):

subtract 55.5 (done on both sides of the equation) divide by 1.65 (done on both sides of the equation)

OpenStudy (aum):

yw.